OpenStudy (anonymous):

Please help me. Solve for m. -3 + m 9 = 10 A. -30 B. 63 C. 87 D. 93. Please help me?

OpenStudy (anonymous):

I am more than happy to help, but the problem is a little confusing, could you rewrite it?

OpenStudy (camerondoherty):

lol ikr @cmorse

OpenStudy (anonymous):

its -3+m / 9 = 10

OpenStudy (camerondoherty):

\[\frac{ -3+m }{ 9 }=10\] Like this?

OpenStudy (anonymous):

exacly

OpenStudy (anonymous):

multiply both sides by 9 9(-3+x)/9=10*9 to get -3+x=90 then add 3 to both sides

OpenStudy (camerondoherty):

its D but let cmorse explain it to you

OpenStudy (anonymous):

thank you both of you <3

OpenStudy (anonymous):

your welcome